how to solve equation
3y+20=3+2y

thank you

Found 2 solutions by Alan3354, oberobic:
Answer by Alan3354(69443)

Subtract 20
then subtract 2y

Answer by oberobic(2304)

3y + 20 = 3 + 2y
Subtract 2y from both sides.
y + 20 = 3
Subtract 20 from both sides.
y = -17
.
Check your work!
3y+20 = -51+20 = -31
3 + 2y = 3 +(-34) = -31
True
.
Done